5. Opportunity cost and production possibilities Eric is a skilled toy maker who is able to produce both boats and drums. He has 8 hours a day to produce toys. The following table shows the daily output resulting from various possible combinations of his time. Hours Producing Produced Choice (Boats) (Drums) (Boats) (Drums) 4 10 4 4 16 18 19
On the following graph, use the blue points (circle symbol) to plot Eric's initial production possibilities frontier (PPF) 30 T 25 Initial PPF 20 New PPF 15 10 4 BOATS
From the previous analysis, you can determine that as Eric increases his production of boats, his opportunity cost of producing one more boat decreases new tool that enables him to produce twice as many boats per hour as before, but it doesn't affect his ability to produce drums (triangle symbol) to plot his new PPF on the previous graph. remains constant Because ne can now make more boats per hour, Eric's opportunity cost of producing drums is it was previously. Grade It Now Save & Continue Continue without saving
Now, suppose Eric is currently using combination C, producing two boats per day. His opportunity cost of producing a third boat per day is per day um us analysis, you can determine that as Eric increases his production of boats, his opportunity cost of producing one more boat 6 drums 10 drums buys a new tool that enables him to produce twice as many boats per hour as before, but it doesn't affect his ability to produce drums. 16 drums points (triangle symbol) to plot his new PPF on the previous graph. Because he can now make more boats per hour, Eric's opportunity cost of producing drums is it was previously Grade It Now Save & Continue Continue without saving
Answer #1

cause he can now make more boats per hour, Erics opportunity cost of producing drums is it was previously
higher than, as if he does not produce drums, he can produce twice the number of boats and opportunity cost of drums is measured in terms of boat

Now, suppose Eric is currently using combination C, producing two boats per day. His opportunity cost of producing a third boHours Producing Produced Choice (Boats) (Drums) (Boats) (Drums) 8 4 10 16

to increase the boats from 2 to 3, drums are reduced from 16 to 10, therefore the opportunity cost is 6

From the previous analysis, you can determine that as Eric increases his production of boats, his opportunity cost of produci
Hours Producing Produced Choice (Boats) (Drums) (Boats) (Drums) 10 6 16 18 2 6 8 19
from 0 to 1, opportunity cost= 19-18 =1
from 1 to 2, opportunity cost= 18-16 =2
and so on
opportunity cost increases

20 0, 19 1,18 2, 16 14 3, 10 S 10 4, 0 0 0.5 1 1.5 22.53 3.5 4 4.5 Boats